# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

### v1

@@ 3 Maurey Sparsification with Functional Value Approximation
$\mathbb{E}\left[F\left(\sum_{t=1}^{T}\lambda_{t}a^{t}\right)-F\left(\sum_{t=1}^{T}\lambda_{t}x^{t}\right)\right]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right].$
Proof.
Let $\bar{x}^{T}=\sum_{t=1}^{T}\lambda_{t}x^{t}$ and $\bar{a}^{T}=\sum_{t=1}^{T}\lambda_{t}a^{t}$. Noting that $F(\bar{a}^{T})=F(\bar{x}^{T}+\bar{a}^{T}-\bar{x}^{T})$, and using repeatedly the $L_{1}$-smoothness of $F$, we have
$F(\bar{a}^{T})$ $\leq F(\bar{x}^{T}+\bar{a}^{T-1}-\bar{x}^{T-1})+\langle\nabla F(\bar{x}^{T}+\bar{a}^{T-1}-\bar{x}^{T-1}),\lambda_{T}(a^{T}-x^{T})\rangle+\frac{L_{1}\lambda_{T}^{2}\left\|a^{T}-x^{T}\right\|^{2}}{2}$
$\ldots$ $\leq F\left(\bar{x}^{T}\right)+\sum_{t\in[T]}\underbrace{\Big{\langle}\nabla F\Big{(}\bar{x}^{T}+\bar{a}^{t-1}-\bar{x}^{t-1}\Big{)},\lambda_{t}(a^{t}-x^{t})\Big{\rangle}}_{\alpha_{t}}+\frac{L_{1}}{2}\sum_{t\in[T]}\lambda_{t}^{2}\|a^{t}-x^{t}\|^{2}$
where $\alpha_{t}=\Big{\langle}\nabla F\Big{(}\sum_{k=t}^{T}\lambda_{k}x^{k}+\sum_{k=1}^{t-1}\lambda_{k}a^{k}\Big{)},\lambda_{t}(a^{t}-x^{t})\Big{\rangle}$ is a martingale difference sequence. Hence, $\mathbb{E}\big{[}\sum_{t\in[T]}\alpha_{t}\big{]}=0$ and $\mathbb{E}[F(\bar{a}^{T})-F(\bar{x}^{T})]\leq\frac{L_{1}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}\mathbb{E}\left[\left\|a^{t}-x^{t}\right\|^{2}\right]$. ∎
We now instantiate Lemma 3.1 for the space $(\mathbb{R}^{d},\|\cdot\|_{1})$ with $\lambda_{i}=\frac{1}{T}$, $i\in[T]$. Here, $x^{t}\in\Delta_{d}$ represents the $t$-th iterate of our algorithm and $a^{t}\sim P_{x^{t}}$ a private, unbiased estimator of $x^{t}$.
Let $F:\mathbb{R}^{d}\to\mathbb{R}$ be a $L_{1}$-smooth function w.r.t. $\|\cdot\|_{1}$. Let $x^{1},...,x^{T}$ be a collection of deterministic vectors in $\Delta_{d}$ and $a^{1},...,a^{T}\in\Delta_{d}$ be random vectors such that $\mathbb{E}[a^{t}|a^{t-1},..,a^{1}]=x^{t}$ for all $t\in[T]$. Then
$\left|\mathbb{E}\left[F\left(\frac{1}{T}\sum_{t=1}^{T}a^{t}\right)-F\left(\frac{1}{T}\sum_{t=1}^{T}x^{t}\right)\right]\right|\leq\frac{2L_{1}}{T}.$
The above result bounds the bias incurred by evaluating gradients at sampled vertices. More concretely, $\left\|\mathbb{E}\left[\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t=1}^{T}x^{t}\right)\right]\right\|_{\infty}\lesssim\frac{L_{2}}{T}$ for second-order-smooth functions since the same bound is true for the bias of each partial derivative (see Corollary D.1). We can also bound the bias for first-order-smooth functions by $\frac{L_{1}}{\sqrt{T}}$, which we do by first approximating the partial derivatives using finite differences, and then using the above corollary directly for the smooth objective function (see Corollary D.2). These results are new, to the best of our knowledge.
We will also need to bound the second-moment of the error incurred by evaluating gradients at sampled vertices. Towards this, first we note that the in-expectation guarantees obtained in Lemma 3.1 can be turned into high-probability guarantees by controlling the increments of the martingale difference sequence $(\alpha_{t})_{t\in[T]}$ with the bounded differences inequality.
Lemma 3.2 (Function value approximation with high probability).
[…]
$\mathbb{P}\left[F\left(\sum_{t=1}^{T}\lambda_{t}a^{t}\right)-F\left(\sum_{t=1}^{T}\lambda_{t}x^{t}\right)\geq\frac{L_{1}D^{2}}{2}\sum_{t=1}^{T}\lambda_{t}^{2}+\beta\frac{L_{0}D}{\sqrt{2}}\sqrt{\sum_{t=1}^{T}\lambda_{t}^{2}}\,\right]\leq\exp(-\beta^{2}).$
We use this high-probability guarantee to give bounds on the expected second-moment error, with respect to $\|\cdot\|_{\infty}$. We have the following result, which we instantiate for two particular cases.
Lemma 3.3 (Second moment of maximum expected error).
Let $F_{1},...,F_{M}:\mathbb{R}^{d}\to\mathbb{R}$ be a collection of $L_{0}$-Lipschitz and $L_{1}$-smooth functions. Let $x^{1},\ldots,x^{T}\in\mathbb{R}^{d}$, $(\lambda_{1},\ldots,\lambda_{T})\in\Delta_{T}$ and $a^{1},...,a^{T}$ be a collection of random vectors such that $\mathbb{E}[a^{t}|a^{t-1},\ldots,a^{1}]=x^{t}$ and $\|a^{t}-x^{t}\|_{1}\leq D$ a.s. for all $t\in[T]$. Let $\bar{a}^{T}=\sum_{t\in[T]}\lambda_{t}a^{t}$ and $\bar{x}^{T}=\sum_{t\in[T]}\lambda_{t}x^{t}$. Then
$\mathbb{E}\left[\max_{j\in[M]}|F_{j}(\bar{x})-F_{j}(\bar{a}^{T})|^{2}\right]\leq\frac{5L_{1}^{2}D^{4}}{4}\left(\sum_{t=1}^{T}\lambda_{t}^{2}\right)^{2}+\frac{L_{0}^{2}D^{2}(6+\log(M))}{2}\sum_{t=1}^{T}\lambda_{t}^{2}.$
Note that when $F$ is second-order-smooth, applying Lemma 3.3 with $F_{j}=\nabla_{j}F$, $j\in[d]$, leads to $\mathbb{E}\|\nabla F\left(\frac{1}{T}\sum_{t\in[T]}a^{t}\right)-\nabla F\left(\frac{1}{T}\sum_{t\in[T]}x^{t}\right)\|^{2}_{\infty}\lesssim\frac{L_{2}^{2}}{T^{2}}+\frac{L_{1}^{2}(1+\log(d))}{T}$. We can also get a bound of $\frac{L_{1}^{2}}{T^{3/2}}+\frac{(L_{0}^{2}+L_{1}^{2})\sqrt{\log(d)}}{\sqrt{T}}$ for first-order-smooth functions by an approximation with finite differences and Lemma 3.3 with $F_{j}(\cdot)=F(\cdot+re_{j})$ for $j\in[d]$. See Corollaries D.3 and  D.4 for details.
4 Differentially Private Stochastic Saddle Points
Problem set-up.
[…]
$\min_{x\in\Delta_{x}}\max_{y\in\Delta_{y}}F_{\mathcal{D}}(x,y),$
under the restriction of using $(\varepsilon,\delta)$-DP algorithms. The efficiency of a candidate solution $(x,y)\in\mathcal{X}\times\mathcal{Y}$, is measured by the duality gap, $\operatorname{Gap}(x,y)=\max_{v\in\mathcal{X},w\in\mathcal{Y}}(F_{\mathcal{D}}(x,w)-F_{\mathcal{D}}(v,y))$: this is a natural measure as it bounds the gap from Nash equilibria when $x,y$ are seen as players in a zero-sum game. Instrumental to the accuracy analysis of algorithms is the so-called saddle operator (a.k.a. the monotone operator associated to the saddle-point problem): given $f:{\cal X}\times{\cal Y}\mapsto\mathbb{R}$, we let $\Phi(x,y)=\big{(}\nabla_{x}f(x,y),-\nabla_{y}f(x,y)\big{)}$ (Rockafellar, 1970).
4.1 Private Stochastic Mirror Descent with Sampling from Vertices
Our first algorithm is a (entropic) stochastic mirror descent (SMD) method with sampling from vertices; see Algorithm 1. A known observation is that sampling from vertices with exponential weights (as induced by SMD) guarantees privacy as it is an instance of the exponential mechanism (e.g., Hsu et al. (2013)). In the bilinear case, this is enough to obtain guarantees for the duality gap as this sampling provides an unbiased first-order stochastic oracle. However, for nonlinear objectives (e.g., general convex-concave losses), it is unclear the accuracy of sampling from vertices.
A key challenge is that the provided extensions of Maurey’s Lemma lead to biased updates. We can reduce this bias to any desired accuracy by increasing the sample size; however, an inherent tradeoff arises: on the one hand, accuracy improves; on the other hand, the privacy degrades. In the following results we optimally tradeoff both considerations under different assumptions on the objective function. Theorem 4.1 considers second-order-smoothness, while Theorem 4.2 only assumes first-order smoothness. We defer the explicit choice of parameters and proof to Appendix E.
Algorithm 1 Private stochastic mirror descent with sequential Maurey sparsification
0:  Dataset $S=\{z^{1},...,z^{n}\}$, Number of steps $T$, Step-size $\tau$, Number of iterate samples $K$
1:  $x^{1}=(1/d_{x},...,1/d_{x})$, $y^{1}=(1/d_{y},...,1/d_{y})$
@@ 4.1 Private Stochastic Mirror Descent with Sampling from Vertices
3:     Let $B^{t}$ be a batch of $B=n/T$ fresh samples
[…]
@@
8:  return  $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$, where $\tilde{x}^{t}\sim P_{x^{t}}$, $\tilde{y}^{t}\sim P_{y^{t}}$
Theorem 4.1 (Guarantees for Algorithm 1 with second order smoothness.).
Suppose the convex-concave objective $f(\cdot,\cdot;z)$ is second-order-smooth. Then there exists $T,\tau$, and $K$ such that Algorithm 1 returns $(\tilde{x},\tilde{y})$ with
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1}+L_{2}){\small\bigg{[}\sqrt{\frac{\log(d_{x})+\log(d_{y})}{n}}+\left(\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{2/5}\bigg{]}.}$
Moreover, if $L_{2}=0$, then there exists $T,\tau$ and $K$ such that Algorithm 1 returns $(\tilde{x},\tilde{y})$ with
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\bigg{[}\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{n}}+\left(\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg{]}.$
In both cases, the output $(\tilde{x},\tilde{y})$ is $(\varepsilon,\delta)$-DP.
@@
The duality gap bound for $L_{2}=0$ (i.e. quadratic objective) in Theorem 4.1 is optimal up to constant factors, at least in the case where $\log(d_{x})\simeq\log(d_{y})$. This follows from a lower bound for the problem of DP Synthetic Data Generation (details in Appendix G). We also note that the proof of this rate only requires $L_{2}$ to be ‘small’ (Remark E.1 for details), rather than exactly 00, so the optimal rates hold for more general nearly quadratic objectives.
Theorem 4.2 (Guarantees for Algorithm 1 with first order smoothness.).
Suppose the convex-concave objective $f(\cdot,\cdot;z)$ is first-order-smooth. Then there exists $T,\tau$ and $K$ such that Algorithm 1 returns an $(\varepsilon,\delta)$-DP point, $(\tilde{x},\tilde{y})$ with
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\left[\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{n}}+\bigg{(}\frac{(\log(d_{x})+\log(d_{y}))^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/3}\bigg{]}.$
To prove these theorems, we will need some technical intermediate results (Appendix E.1 for proofs). The first is a condition on the stepsize that ensures the output $(\tilde{x},\tilde{y})$ is private.
@@
Algorithm 1 is $(\varepsilon,\delta)$-DP if $\tau\leq B\varepsilon/[8L_{0}\sqrt{2T(K+1)\log(1/\delta)}].$
@@
For the accuracy analysis, note that Algorithm 1 produces both a private sequence $(\tilde{x}^{t},\tilde{y}^{t})_{t\in[T]}$ and a non-private one $(x^{t},y^{t})_{t\in[T]}$. We first show a bound for the average of the non-private iterates.
Lemma 4.1 (Implicit in (Nemirovski et al., 2009)).
@@
Let $(x^{t},y^{t})$ be the (non-private) iterates of Algorithm 1 generated in line 6, and let $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t}$ and $\bar{y}=\sum_{t\in[T]}y^{t}$. Furthermore, let $g^{t}_{x}$ and $g^{t}_{y}$ be the estimators constructed in line 5 of the algorithm. Then, the following inequality holds.
$T\operatorname{Gap}(\bar{x},\bar{y})$ $\leq\frac{2(\log(d_{x})+\log(d_{y}))}{\tau}+\tau\sum_{t=1}^{T}(\|g^{t}\|_{\infty}^{2}+\|\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\|_{\infty}^{2})$
$\qquad+\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
where $g^{t}=(g^{t}_{x},g^{t}_{y})$ and the sequences $(w^{t})_{t\in[T]}$ and $(v^{t})_{t\in[T]}$ are completely determined by the randomness of the algorithm up to iteration $t$.
@@
For explicit definitions of $(w^{t},v^{t})$ and the proof of Lemma 4.1, see Appendix E.1. Now we quantify the discrepancy between the gap of the private (sampled) and non-private iterates. Since the duality gap is defined in terms of function values, we would like to use Maurey’s Sparsification Lemma. However, since the duality gap involves a maximum –and the maximum of smooth functions is not necessarily smooth– the sparsification lemma does not directly apply. We handle this issue with a smoothing argument, and get the following lemma (see Appendix E.1 for proof).

### v2

@@ 3.1 Private Stochastic Mirror Descent with Sampling from Vertices
3:  $(\hat{x}^{t},\hat{y}^{t})=\frac{1}{K}\sum_{k\in[K]}(\hat{x}^{t,k},\hat{y}^{t,k})$, where $(\hat{x}^{t,k},\hat{y}^{t,k})_{k\in[K]}\overset{\text{iid}}{\sim}P_{x^{t}}\times P_{y^{t}}$
4:  Let $B^{t}$ be a batch of $B=n/T$ fresh samples
@@
8:return $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$, where $(\tilde{x}^{t},\tilde{y}^{t})_{t\in[T]}\overset{\text{iid}}{\sim}P_{x^{t}}\times P_{y^{t}}$
Theorem 3.1 (Guarantees for Algorithm 1).
[…]
(5) $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\left[\sqrt{\frac{\ell}{n}}+\bigg(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/3}\bigg].$
If in addition $f(\cdot,\cdot;z)$ has $L_{2}$-smooth partial derivatives, then there exists a different setting for $T,\tau$ and $K$ such that Algorithm 1 is $(\varepsilon,\delta)$-DP and its output satisfies
(6) $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1}+L_{2})\bigg[\sqrt{\frac{\ell}{n}}+\left(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{2/5}\bigg].$
[…]
(7) $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim(L_{0}+L_{1})\bigg[\sqrt{\frac{\ell}{n}}+\left(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg].$
@@
The duality gap bound for $L_{2}=0$ (i.e. quadratic objective) in Theorem 3.1 is optimal up to constant factors in the case where $\log(d_{x})\simeq\log(d_{y})$. This follows from the lower bound in [lower_bound_DP_query_answering] with a resampling argument from [Bassily:2019]. Moreover, the proof of this rate only requires $L_{2}\lesssim(L_{0}+L_{1})\Big\{\sqrt{\ell/n}+\sqrt{\ell^{3/2}\sqrt{\log(1/\delta)}/(n\varepsilon)}\Big\}$, so the optimal rates hold for more general nearly quadratic objectives.
3.1.1 Privacy Analysis of Algorithm 1
We prove that the algorithm is DP. More concretely, we provide a condition on the stepsize that suffices to get privacy.
@@ 3.1.1 Privacy Analysis of Algorithm 1
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. Then, Algorithm 1 is $(\varepsilon,\delta)$-DP if $\tau\leq\frac{B\varepsilon}{16L_{0}\sqrt{T(K+1)\log(1/\delta)}}$.
The proof relies on the privacy guarantees of the exponential mechanism and on the advanced composition property of DP. For completeness, we include them below.
Theorem 3.4 (Exponential mechanism (Theorem 3.10 in [dwork2014DPfoundations])).
Consider a (score) function $s:{\cal Z}^{n}\times J\mapsto\mathbb{R}$, where $J=\{e_{i}\}_{i\in[d]}$. Define the sensitivity of $s$ as $\Delta_{s}:=\max_{e\in J}\max_{S,S^{\prime}\in\mathcal{Z}^{n}\text{ differing in }\leq 1\text{ element}}|s(S,e)-s(S^{\prime},e)|$. Then, sampling an object $e$ from $J$ according to $\mathbb{P}(e=e_{i})\propto\exp\left(\frac{\varepsilon s(S,e_{i})}{2\Delta_{s}}\right)$ is $\varepsilon$-DP. As a consequence, sampling $e$ according to $\mathbb{P}(e=e_{i})\propto\exp\left(s(S,e_{i})\right)$ is $2\Delta_{s}$-DP.
Theorem 3.5 (Fully Adaptive Composition [Whitehouse:2023]).
Suppose $(\mathcal{A}_{n})_{n\geq 1}$ is a sequence of algorithms such that, for any $n\geq 1$, $\mathcal{A}_{n}$ is $(\varepsilon_{n},\delta_{n})$-DP with respect to a dataset $S$ when conditioned on $\mathcal{A}_{1:n-1}$. Let $\varepsilon\geq 0$ and $\delta=\delta^{\prime}+\delta^{\prime\prime}$ be the target privacy parameters such that $\delta^{\prime},\delta^{\prime\prime}>0$. Consider the function $N:\mathbb{R}^{\infty}_{\geq 0}\times\mathbb{R}^{\infty}_{\geq 0}\to\mathbb{N}$ such that $N((\varepsilon_{n})_{n\geq 1},(\delta_{n})_{n\geq 1})$ is defined as
$\inf\left\{n:\varepsilon<\sqrt{2\log\left(\frac{1}{\delta^{\prime}}\right)\sum_{m\leq n+1}\varepsilon_{m}^{2}}+\frac{1}{2}\sum_{\sum_{m\leq n+1}}\varepsilon_{m}^{2}\text{ or }\delta^{\prime\prime}<\sum_{m\leq n+1}\delta_{m}\right\}.$
The following corollary follows from the theorem above, by noting that $\varepsilon^{\prime}\leq\frac{\varepsilon}{2\sqrt{2T\log(1/\delta)}}$ and $0\leq\varepsilon\leq 8\log(1/\delta)$ imply $\varepsilon\geq\varepsilon^{\prime 2}T/2+\varepsilon^{\prime}\sqrt{2T\log(1/\delta)}$.
Corollary 3.6 (Advanced composition of pure DP mechanisms).
For $\varepsilon^{\prime}>0$. the $T$-fold adaptive composition of $\varepsilon^{\prime}$-DP mechanisms satisfies $(\epsilon,\delta)$-DP for any $\delta\in(0,1)$ as long as $\varepsilon^{\prime}\leq\frac{\varepsilon}{2\sqrt{2T\log(1/\delta)}}$ and $0\leq\varepsilon\leq 8\log(1/\delta)$.
Proof 3.7 (Proof of Proposition 3.3).
We start proving that for all $t\in[T],k\in[K]$, $\tilde{x}^{t+1},\tilde{y}^{t+1}$ and $\hat{x}^{t+1,k},\hat{y}^{t+1,k}$ are private given the previous private iterates. First, notice that $\tilde{x}^{1},\tilde{y}^{1}$ and $\hat{x}^{1,k},\hat{y}^{1,k}$ are trivially private because they don’t depend on the dataset $S$. Now we consider $\tilde{x}^{t+1},\hat{x}^{t+1,k}$. Unrolling the $x$ update, $x^{t+1}_{j}\propto\exp\left\{-\tau\left(\sum_{i=1}^{t}g^{i}_{x,j}\right)\right\}$, where $g^{t}_{x,j}$ denotes the $j$-th coordinate of $g^{t}_{x}$ for every $t\in[T]$. Hence, $\tilde{x}^{t+1}$ and $\hat{x}^{t+1,k}$ are vertices of the simplex $\Delta_{x}$, and each of them is choosen with the exponential mechanism with score function $s^{t}_{x}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{x,j}\right)$. The privacy guarantee of the exponential mechanism in Theorem 3.4 implies that $\tilde{x}^{t+1}$ and $\hat{x}^{t+1,k}$ are $2\Delta(s^{t}_{x})$-DP, where $\Delta(s^{t}_{x})$ is the sensitivity of $s^{t}_{x}$. Using the same idea, we argue that $\tilde{y}^{t+1}$ and $\hat{y}^{t+1,k}$ are all $2\Delta(s^{t}_{y})$-DP, where $\Delta(s^{t}_{y})$ is the sensitivity of $s^{t}_{y}(S,j)=-\tau\left(\sum_{i=1}^{t}g^{i}_{y,j}\right)$.
By $L_{0}$-Lipschitzness of $f$ we have that $\max_{j\in[d]}\left|s^{t}_{x}(S,j)-s^{t}_{x}(S^{\prime},j)\right|\leq\frac{2L_{0}\tau}{B}$. Hence, every sampled vertex $\tilde{x}^{t+1},\hat{x}^{t+1,k}$ is $\frac{4L_{0}\tau}{B}$-DP, conditionally on the previously sampled vertices. Analogously, $\tilde{y}^{t+1}$ and $\hat{y}^{t+1,k}$ are $\frac{4L_{0}\tau}{B}$-DP, conditionally the on previously sampled vertices. By Advanced Composition of DP (Corollary 3.6), it suffices that each sampled vertex is $\frac{\varepsilon}{2\sqrt{4T(K+1)\log(1/\delta)}}$-DP in order to make the entire Algorithm 1 $(\varepsilon,\delta)$-DP. Then, for the algorithm to be $(\varepsilon,\delta)$-DP, it suffices that $\frac{4L_{0}\tau}{B}\leq\frac{\varepsilon}{2\sqrt{4T(K+1)\log(1/\delta)}}$.
3.1.2 Convergence Analysis of Algorithm 1
@@
For the accuracy analysis, note that Algorithm 1 produces both a private sequence $(\tilde{x}^{t},\tilde{y}^{t})_{t\in[T]}$ and a non-private one $(x^{t},y^{t})_{t\in[T]}$. We will bound the expected duality gap evaluated at the private iterates’ average. To do this, we first show a gap bound on the non-private iterates’ average, and then we quantify the discrepancy between the duality gap evaluated at the private and non-private iterates’ average.
Lemma 3.8 (Duality Gap at non-private iterates’ average (Implicit in [JNLS:2009])).
@@
Let $(x^{t},y^{t})$ be the (non-private) iterates of Algorithm 1 generated in line 6, and let $\bar{x}=\frac{1}{T}\sum_{t\in[T]}x^{t}$ and $\bar{y}=\frac{1}{T}\sum_{t\in[T]}y^{t}$. Furthermore, let $g^{t}_{x}$ and $g^{t}_{y}$ be the estimators constructed in line 5 of the algorithm and $g^{t}=(g^{t}_{x},g^{t}_{y})$. Denote $\widetilde{\Delta}_{t}=\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}$. Then, for some sequences $(w^{t})_{t\in[T]}\subset\Delta_{x}$ and $(v^{t})_{t\in[T]}\subset\Delta_{y}$ that are predictable w.r.t. the natural filtration, the following inequality holds
$T\operatorname{Gap}(\bar{x},\bar{y})$ $\leq\frac{2\ell}{\tau}+\tau\sum_{t=1}^{T}(\|g^{t}\|_{\infty}^{2}+\|\widetilde{\Delta}_{t}\|_{\infty}^{2})+\sum_{t=1}^{T}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
For the proof of this lemma, we will use the following auxiliary result. This result is well known in the SSP literature, and follows directly from regret bounds for the (online) mirror-descent method. Similarly to other works, we use this result in order to circumvent uniform convergence to bound the gap function.
Lemma 3.9 (Corollary 2 in [juditsky_VI]).
Let $\xi^{1},\xi^{2},$... be a sequence in $\mathbb{R}^{d}$ and $\mathcal{X}$ be a compact, convex set. Let $\psi:\mathbb{R}^{d}\to\mathbb{R}$ be a $1$-strongly convex w.r.t $\|\cdot\|$ and $D_{\psi}(x,y)=\psi(x)-\psi(y)-\langle\nabla\psi(y),x-y\rangle$. Define $w^{1}\in\mathcal{X}$, $w^{t+1}=\operatorname{argmin}_{x\in\mathcal{X}}D_{\psi}(w^{t},x)+\langle\xi^{t},x\rangle$. Then, for all $w\in\mathcal{X}$ and $T\in\mathbb{N}$: $\sum_{t\in[T]}\langle\xi^{t},w^{t}-w\rangle\leq D_{\psi}(w^{1},w)+\frac{1}{2}\sum_{t\in[T]}\|\xi^{t}\|^{2}_{*}$, where $\|\cdot\|_{*}$ is the dual norm of $\|\cdot\|$.
Proof 3.10 (Proof of Lemma 3.8).
By convexity-concavity of the loss and the convergence rate of Stochastic Mirror Descent in the simplex setup [JNLS:2009, juditsky_VI]:
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]\leq\sum_{t\in[T]}F_{\mathcal{D}}(x^{t},v)-F_{\mathcal{D}}(w,y^{t})$
$\leq\sum_{t\in[T]}\langle\Phi_{\mathcal{D}}(x^{t},y^{t}),(x^{t},y^{t})-(w,v)\rangle\leq\frac{\ell}{\tau}+\tau\sum_{t\in[T]}\|g^{t}\|_{\infty}^{2}+\sum_{t=1}^{T}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w,v)\rangle,$
for all $w\in\Delta_{x},v\in\Delta_{y}$. Lemma 3.9 with $\xi^{t}=\tau(\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x})$, $\psi$ the entropy function, $\|\cdot\|=\|\cdot\|_{1}$ and $w^{1}=(1/d_{x},...,1/d_{x})$ gives
Using this result and denoting $\widetilde{\Delta}_{t,x}=\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}$ we can conclude that
$\sum_{t\in[T]}\langle\widetilde{\Delta}_{t,x},x^{t}-w\rangle\leq\sum_{t\in[T]}\langle\widetilde{\Delta}_{t,x},x^{t}-w^{t}\rangle+\frac{\log(d_{x})}{\tau}+\frac{\tau}{2}\sum_{t\in[T]}\|\nabla_{x}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x}\|_{\infty}^{2}.$
An analogous bound for the $y$ iterate holds. Hence, for any $w\in\Delta_{x},v\in\Delta_{y}$,
$T[F_{\mathcal{D}}(\bar{x},v)-F_{\mathcal{D}}(w,\bar{y})]$ $\leq\frac{2\ell}{\tau}+\tau\sum_{t=1}^{T}(\|g^{t}\|_{\infty}^{2}+\|\widetilde{\Delta}_{t}\|_{\infty}^{2})+\sum_{t=1}^{T}\langle\widetilde{\Delta}_{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
Maximizing over $w\in\Delta_{x},v\in\Delta_{y}$ finishes the proof.
@@
Now we quantify the discrepancy between the gap of the private (sampled) and non-private iterates. The duality gap is defined in terms of function values, but since it involves a maximum (and the maximum of smooth functions is not necessarily smooth) our sparsification results do not directly apply. We handle this issue with a smoothing argument that leads to the following lemma.